OpenStudy (anonymous):

A car travels along a straight line at a constant speed of 60.5 mi/h for a distance d and then another distance d in the same direction at another constant speed. The average velocity for the entire trip is 27.0 mi/h. (a) What is the constant speed with which the car moved during the second distance d?

OpenStudy (fifciol):

average velocity= total distance/ total time=2d/total time

OpenStudy (fifciol):

total time is t1 + t2 t1=d/60.5 t2=d/v substitute and solve for v

OpenStudy (fifciol):

\[27=\frac{ 2d }{ t_1+t_2 }=\frac{ 2d }{ \frac{ d }{ 60.5 }+\frac{ d }{ v } }=\frac{ 2d }{ \frac{ dv+60.5d }{ 60.5v } }=\frac{ 2d*60.5v }{d(v+60.5) }=\frac{ 121v }{ v+60.6 }\]

OpenStudy (fifciol):

27(v+60.5)=121v

OpenStudy (fifciol):

v=17.38 mi/h

OpenStudy (anonymous):

I absolutely can't thank you enough-- at this very moment i'm at a some starbucks, i've been beating my head against the wall over this problem. I've been through considerable amounts of math, but apparently coming into physics with a boastful attitude has proven to be of great handicap on my part. I thank you again!